# Asymptotic theory of in-context learning by linear attention

### v2

@@ Acknowledgments
Next, we replace the “leave-one-out” terms $G_{e}^{\mu}$ and $\chi^{\mu}(w_{\mu})$ in (SI.4.21) by their “full” versions. Specifically, we replace $G_{e}^{\mu}$ by $G_{e}$, and $\chi^{\mu}(w_{\mu})$ by
$\chi(w_{\mu})\coloneqq\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{\mu})\right]\right).$ (SI.4.22)
It is important to note the difference between (SI.4.13) and (SI.4.22): the former uses $G_{e}^{\mu}$ and the latter $G_{e}$. After these replacements and using $\Delta_{4}$ to capture the approximation errors, we have
$G_{e}\left(\frac{\tau}{n}\sum_{\mu\in[n]}\frac{1}{1+\chi(w_{\mu})}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)=I+\sum_{j\leq 4}\Delta_{j}.$ (SI.4.23)
Recall that there are $k$ unique task vectors $\left\{w_{i}\right\}_{1\leq i\leq k}$ in the training set consisting of $n$ input samples. Each sample is associated with one of these task vectors, sampled uniformly from the set $\left\{w_{i}\right\}_{1\leq i\leq k}$. In our analysis, we shall assume that $k$ divides $n$ and that each unique task vector is associated with exactly $n/k$ input samples. (We note that this assumption merely serves to simplify the notation. The asymptotic characterization of the random matrix $G_{e}$ remains the same even without this assumption.) Observe that there are only $k$ unique terms in the sum on the left-hand side of (SI.4.23). Thus,
$G_{e}\left(\frac{\tau}{k}\sum_{i\in[k]}\frac{1}{1+\chi(w_{i})}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{i}^{\top}\otimes\begin{bmatrix}w_{i}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{i}\otimes\begin{bmatrix}w_{i}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{i})\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)=I+\sum_{j\leq 4}\Delta_{j}.$ (SI.4.24)
So far, we have been treating the $k$ task vectors $\left\{w_{i}\right\}_{i\in[k]}$ as fixed vectors, only using the randomness in the input samples that are associated with the data vectors $\left\{x_{i}^{\mu}\right\}$. To further simplify our asymptotic characterization, we take advantage of the fact that $\left\{w_{i}\right\}_{i\in[k]}$ are independently sampled from $\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$. To that end, we can first show that $\chi(w_{i})$ in (SI.4.22) concentrates around its expectation. Specifically,
$\chi(w_{i})=\mathbb{E}\left[\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{i})\right]\right)\right]+\mathcal{O}_{\prec}(d^{-1/2}).$ (SI.4.25)
@@
$\mathbb{E}\left[\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{i})\right]\right)\right]=\mathbb{E}\left[\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{j})\right]\right)\right]$ (SI.4.26)
for any $1\leq i<j\leq k$. It follows that $\lvert\chi(w_{i})-\chi(w_{j})\rvert=\mathcal{O}_{\prec}(d^{-1/2})$, and thus, by a union bound,
$\max_{i\in[k]}\,\lvert\chi(w_{k_{1}})-\widehat{\chi}_{\mathrm{ave}}\rvert=\mathcal{O}_{\prec}(d^{-1/2}),$ (SI.4.27)
@@
$\widehat{\chi}_{\mathrm{ave}}\coloneqq\frac{1}{k}\sum_{i\in[k]}\chi(w_{i}).$ (SI.4.28)
Upon substituting (SI.4.22) into (SI.4.28), it is straightforward to verify the following characterization of $\widehat{\chi}_{\mathrm{ave}}$:
$\widehat{\chi}_{\mathrm{ave}}=\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right).$ (SI.4.29)
The estimate in (SI.4.27) prompts us to replace the terms $\chi(w_{i})$ in the right-hand side of (SI.4.24) by the common value $\widehat{\chi}_{\mathrm{ave}}$. As before, we introduce a matrix $\Delta_{5}$ to capture the approximation error associated with this step. Using the newly introduced notation $E_{\mathrm{tr}},b_{\mathrm{tr}}$ and $R_{\mathrm{tr}}$ in (SI.4.3) and (SI.4.1), we can then simplify (SI.4.24) as
$G_{e}\left(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)$ (SI.4.30)
@@
$\widehat{\mathcal{G}}_{e}(\pi)\coloneqq\left(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)^{-1}.$ (SI.4.31)
@@
$G_{e}=\widehat{\mathcal{G}}_{e}(\pi)+\underbrace{\widehat{\mathcal{G}}_{e}(\pi)\left(\Delta_{1}+\Delta_{2}+\Delta_{3}+\Delta_{4}+\Delta_{5}\right)}_{\text{approximation errors}}.$ (SI.4.32)
@@
We claim that $\widehat{\mathcal{G}}_{e}$ is asymptotically equivalent to $G_{e}$, in the sense of Definition 1. Given (SI.4.32), proving this claim requires showing that, for $j=1,2,\ldots,5$,
$u^{\top}\left(\widehat{\mathcal{G}}_{e}(\pi)\Delta_{j}\right)v\simeq 0,$ (SI.4.33a)
$\frac{1}{\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A_{\mathrm{tr}})^{\top}\end{bmatrix}\left(\widehat{\mathcal{G}}_{e}(\pi)\Delta_{j}\right)v\simeq 0,$ (SI.4.33b)
@@
$\frac{1}{d^{2}}\operatorname{tr}\left(\left[\widehat{\mathcal{G}}_{e}(\pi)\Delta_{j}\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right)\simeq 0,$ (SI.4.33c)
@@
We note the equivalent matrix $\widehat{\mathcal{G}}_{e}(\pi)$ still involves one scalar $\widehat{\chi}_{\mathrm{ave}}$ that depends on the original resolvent $G_{e}(\pi)$. Next, we show that $\widehat{\chi}_{\mathrm{ave}}$ can be replaced by $\chi_{\pi}$, the unique positive solution to (SI.4.7). To that end, we recall the characterization in (SI.4.29). Using the claim that $G_{e}(\pi)$ and $\widehat{\mathcal{G}}_{e}(\pi)$ are asymptotically equivalent (in particular, in the sense of (SI.4.6)), we have
$\widehat{\chi}_{\mathrm{ave}}\simeq\frac{1}{d^{2}}\operatorname{tr}\left(\left[\widehat{\mathcal{G}}_{e}(\pi)\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right).$ (SI.4.34)
To compute the first term on the right-hand side of the above estimate, we directly invert the block matrix $\widehat{\mathcal{G}}_{e}(\pi)$ in (SI.4.31). Recall that $\Pi_{e}$ is chosen in the forms of (SI.3.6) and (SI.3.13). It is then straightforward to verify that
$\widehat{\mathcal{G}}_{e}=\begin{bmatrix}\bar{c}&-\bar{c}\,{\bar{q}}^{\top}\\ -\bar{c}\,\bar{q}&I\otimes F_{E}(\widehat{\chi}_{\mathrm{ave}})+\bar{c}\,\bar{q}{\bar{q}}^{\top}\end{bmatrix},$ (SI.4.35)
@@
$F_{E}(\chi)=\Big{(}\frac{\tau}{1+\chi}E_{\mathrm{tr}}+\pi B+\lambda\tau I_{d+1}\Big{)}^{-1},$ (SI.4.36)
$\bar{q}=\frac{\tau}{(1+\widehat{\chi}_{\mathrm{ave}})\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\widehat{\chi}_{\mathrm{ave}})\right),$ (SI.4.37)
and
${1/}\bar{c}=\frac{\tau(1+\rho)}{1+\widehat{\chi}_{\mathrm{ave}}}+\lambda\tau-\frac{\tau^{2}}{(1+\widehat{\chi}_{\mathrm{ave}})^{2}d}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\widehat{\chi}_{\mathrm{ave}})\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}^{\top}\right).$ (SI.4.38)
Using (SI.4.35), we can now write the equation (SI.4.34) as
$\widehat{\chi}_{\mathrm{ave}}$ $\simeq\frac{1}{d}\operatorname{tr}\left(F_{E}(\widehat{\chi}_{\mathrm{ave}})E_{\mathrm{tr}}\right)$ (SI.4.39)
$\quad+\frac{\bar{c}\,\tau^{2}}{(1+\widehat{\chi}_{\mathrm{ave}})^{2}d^{3}}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\widehat{\chi}_{\mathrm{ave}})E_{\mathrm{tr}}F_{E}(\widehat{\chi}_{\mathrm{ave}})\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}^{\top}\right).$
The second term on the right-hand side of (SI.4.39) is negligible. Indeed,
$\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\widehat{\chi}_{\mathrm{ave}})E_{\mathrm{tr}}F_{E}(\widehat{\chi}_{\mathrm{ave}})\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}^{\top}\right)$ (SI.4.40)
$\qquad\qquad\qquad\leq\lVert F_{E}(\widehat{\chi}_{\mathrm{ave}})E_{\mathrm{tr}}F_{E}(\widehat{\chi}_{\mathrm{ave}})\rVert_{\mathsf{op}}(\,\lVert R_{\mathrm{tr}}\rVert_{\mathsf{F}}^{2}+(1+\rho)^{2}\lVert b_{\mathrm{tr}}\rVert^{2}).$
By construction, $\lVert F_{E}(\widehat{\chi}_{\mathrm{ave}})\rVert_{\mathsf{op}}\leq(\lambda\tau)^{-1}$. Moreover, since the task vectors $\left\{w_{i}\right\}_{i\in[k]}$ are independent vectors sampled from $\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$, it is easy to verify that
$\lVert E_{\mathrm{tr}}\rVert_{\mathsf{op}}=\mathcal{O}_{\prec}(1),\qquad\lVert R_{\mathrm{tr}}\rVert_{\mathsf{F}}=\mathcal{O}_{\prec}(\sqrt{d})\qquad\text{and}\qquad\lVert b_{\mathrm{tr}}\rVert_{2}=\mathcal{O}_{\prec}(1).$ (SI.4.41)
Finally, since $\bar{c}$ is an element of $\widehat{\mathcal{G}}_{e}$, we must have $\lvert\bar{c}\rvert\leq\lVert\widehat{\mathcal{G}}_{e}\rVert_{\mathsf{op}}\leq(\tau\lambda)^{-1}$. Combining these estimates gives us
$\frac{\bar{c}\,\tau^{2}}{(1+\widehat{\chi}_{\mathrm{ave}})^{2}d^{3}}\operatorname{tr}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\widehat{\chi}_{\mathrm{ave}})E_{\mathrm{tr}}F_{E}(\widehat{\chi}_{\mathrm{ave}})\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}^{\top}\right)=\mathcal{O}_{\prec}(d^{-2}),$ (SI.4.42)
and thus we can simplify (SI.4.39) as
$\widehat{\chi}_{\mathrm{ave}}\simeq\frac{1}{d}\operatorname{tr}\left[\Big{(}\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}E_{\mathrm{tr}}+\pi B+\lambda\tau I_{d}\Big{)}^{-1}E_{\mathrm{tr}}\right].$ (SI.4.43)
Observe that (SI.4.43) is a small perturbation of the self-consistent equation in (SI.4.7). By the stability of the equation (SI.4.7), we then have
$\widehat{\chi}_{\mathrm{ave}}\simeq\chi_{\pi},$ (SI.4.44)
where $\chi_{\pi}$ is the unique positive solution to (SI.4.7).
Recall the definitions of $\mathcal{G}_{e}(\pi)$ and $\widehat{\mathcal{G}}_{e}(\pi)$ in (SI.4.31) and (SI.4.8), respectively. By the standard resolvent identity,
$\widehat{\mathcal{G}}_{e}(\pi)-\mathcal{G}_{e}(\pi)$ (SI.4.45)
By construction, $\lVert\widehat{\mathcal{G}}_{e}(\pi)\rVert_{\mathsf{op}}\leq 1/(\tau\lambda)$ and $\lVert\mathcal{G}_{e}(\pi)\rVert_{\mathsf{op}}\leq 1/(\tau\lambda)$. Moreover, $\lVert E_{\mathrm{tr}}\rVert_{\mathsf{op}}\prec 1$ and
$\lVert\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)\rVert\prec 1.$ (SI.4.46)
It then follows from (SI.4.44) and (SI.4.45) that
$\lVert\widehat{\mathcal{G}}_{e}(\pi)-\mathcal{G}_{e}(\pi)\rVert_{\mathsf{op}}\simeq 0.$ (SI.4.47)
If $\widehat{\mathcal{G}}_{e}(\pi)$ satisfies the equivalent conditions (SI.4.4), (SI.4.5) and (SI.4.6) (as claimed in our analysis above), then the estimate in (SI.4.47) allows us to easily check that $\mathcal{G}_{e}(\pi)$ also satisfies (SI.4.4), (SI.4.5) and (SI.4.6). Thus, we claim that $\mathcal{G}_{e}(\pi)$ is asymptotically equivalent to the extended resolvent matrix $G_{e}(\pi)$ in the sense of Definition 1.
SI.5 Asymptotic Limits of the Generalization Errors
SI.5.1 Asymptotic Limits of the Linear and Quadratic Terms
From Corollary 1 and the discussions in Remark 2, characterizing the test error $e(\Gamma^{\ast})$ boils down to computing the linear term $\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}A_{\mathrm{test}}^{\top}\right)$ and the quadratic term $\frac{1}{d}\operatorname{tr}\left(\Gamma^{\ast}B_{\mathrm{test}}(\Gamma^{\ast})^{\top}\right)$, where $A_{\mathrm{test}}$ and $B_{\mathrm{test}}$ are the matrices defined in (SI.2.23) and (SI.2.24), respectively.
We consider two different types of test data distributions $\mathcal{P}_{\mathrm{test}}$: ICL and IDG. [See Section II.6 in the main text for details.] From (SI.2.16) and (SI.2.17), these two settings correspond to choosing
[…]

### v3

@@ Appendix D An Asymptotic Equivalent of the Extended Resolvent Matrix
Next, we replace the “leave-one-out” terms $G_{e}^{\mu}$ and $\chi^{\mu}(w_{\mu})$ in (120) by their “full” versions. Specifically, we replace $G_{e}^{\mu}$ by $G_{e}$, and $\chi^{\mu}(w_{\mu})$ by
$\chi(w_{\mu})\coloneqq\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{\mu})\right]\right).$ (121)
It is important to note the difference between (112) and (121): the former uses $G_{e}^{\mu}$ and the latter $G_{e}$. After these replacements and using $\Delta_{4}$ to capture the approximation errors, we have
$G_{e}\left(\frac{\tau}{n}\sum_{\mu\in[n]}\frac{1}{1+\chi(w_{\mu})}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{\mu}^{\top}\otimes\begin{bmatrix}w_{\mu}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{\mu}\otimes\begin{bmatrix}w_{\mu}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{\mu})\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)=I+\sum_{j\leq 4}\Delta_{j}.$ (122)
Recall that there are $k$ unique task vectors $\left\{w_{i}\right\}_{1\leq i\leq k}$ in the training set consisting of $n$ input samples. Each sample is associated with one of these task vectors, sampled uniformly from the set $\left\{w_{i}\right\}_{1\leq i\leq k}$. In our analysis, we shall assume that $k$ divides $n$ and that each unique task vector is associated with exactly $n/k$ input samples. (We note that this assumption merely serves to simplify the notation. The asymptotic characterization of the random matrix $G_{e}$ remains the same even without this assumption.) Observe that there are only $k$ unique terms in the sum on the left-hand side of (122). Thus,
$G_{e}\left(\frac{\tau}{k}\sum_{i\in[k]}\frac{1}{1+\chi(w_{i})}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}w_{i}^{\top}\otimes\begin{bmatrix}w_{i}^{\top}&1+\rho\end{bmatrix}\\ \frac{1}{\sqrt{d}}w_{i}\otimes\begin{bmatrix}w_{i}\\ 1+\rho\end{bmatrix}&I_{d}\otimes E(w_{i})\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)=I+\sum_{j\leq 4}\Delta_{j}.$ (123)
So far, we have been treating the $k$ task vectors $\left\{w_{i}\right\}_{i\in[k]}$ as fixed vectors, only using the randomness in the input samples that are associated with the data vectors $\left\{x_{i}^{\mu}\right\}$. To further simplify our asymptotic characterization, we take advantage of the fact that $\left\{w_{i}\right\}_{i\in[k]}$ are independently sampled from $\mathsf{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$. To that end, we can first show that $\chi(w_{i})$ in (121) concentrates around its expectation. Specifically,
$\chi(w_{i})=\mathbb{E}\left[\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{i})\right]\right)\right]+\mathcal{O}_{\prec}(d^{-1/2}).$ (124)
@@
$\mathbb{E}\left[\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{i})\right]\right)\right]=\mathbb{E}\left[\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E(w_{j})\right]\right)\right]$ (125)
for any $1\leq i<j\leq k$. It follows that $\mathinner{\!\left\lvert\chi(w_{i})-\chi(w_{j})\right\rvert}=\mathcal{O}_{\prec}(d^{-1/2})$, and thus, by a union bound,
$\max_{i\in[k]}\,\mathinner{\!\left\lvert\chi(w_{k_{1}})-\widehat{\chi}_{\mathrm{ave}}\right\rvert}=\mathcal{O}_{\prec}(d^{-1/2}),$ (126)
@@
$\widehat{\chi}_{\mathrm{ave}}\coloneqq\frac{1}{k}\sum_{i\in[k]}\chi(w_{i}).$ (127)
Upon substituting (121) into (127), it is straightforward to verify the following characterization of $\widehat{\chi}_{\mathrm{ave}}$:
$\widehat{\chi}_{\mathrm{ave}}=\frac{1}{d^{2}}\operatorname{tr}\left(\left[G_{e}\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right).$ (128)
The estimate in (126) prompts us to replace the terms $\chi(w_{i})$ in the right-hand side of (123) by the common value $\widehat{\chi}_{\mathrm{ave}}$. As before, we introduce a matrix $\Delta_{5}$ to capture the approximation error associated with this step. Using the newly introduced notation $E_{\mathrm{tr}},b_{\mathrm{tr}}$ and $R_{\mathrm{tr}}$ in (102) and (100), we can then simplify (123) as
$G_{e}\left(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)$ (129)
@@
$\widehat{\mathcal{G}}_{e}(\pi)\coloneqq\left(\frac{\tau}{1+\widehat{\chi}_{\mathrm{ave}}}\begin{bmatrix}1+\rho&\frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)^{\top}\\ \frac{1}{\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}\right)&I_{d}\otimes E_{\mathrm{tr}}\end{bmatrix}+\pi\Pi_{e}+\tau\lambda I\right)^{-1}.$ (130)
@@
$G_{e}=\widehat{\mathcal{G}}_{e}(\pi)+\underbrace{\widehat{\mathcal{G}}_{e}(\pi)\left(\Delta_{1}+\Delta_{2}+\Delta_{3}+\Delta_{4}+\Delta_{5}\right)}_{\text{approximation errors}}.$ (131)
@@
We claim that $\widehat{\mathcal{G}}_{e}$ is asymptotically equivalent to $G_{e}$, in the sense of Definition 1. Given (131), proving this claim requires showing that, for $j=1,2,\ldots,5$,
$u^{\top}\left(\widehat{\mathcal{G}}_{e}(\pi)\Delta_{j}\right)v\simeq 0,$ (132a)
$\frac{1}{\sqrt{d}}\begin{bmatrix}0&\operatorname{vec}(A_{\mathrm{tr}})^{\top}\end{bmatrix}\left(\widehat{\mathcal{G}}_{e}(\pi)\Delta_{j}\right)v\simeq 0,$ (132b)
@@
$\frac{1}{d^{2}}\operatorname{tr}\left(\left[\widehat{\mathcal{G}}_{e}(\pi)\Delta_{j}\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right)\simeq 0,$ (132c)
@@
We note the equivalent matrix $\widehat{\mathcal{G}}_{e}(\pi)$ still involves one scalar $\widehat{\chi}_{\mathrm{ave}}$ that depends on the original resolvent $G_{e}(\pi)$. Next, we show that $\widehat{\chi}_{\mathrm{ave}}$ can be replaced by $\chi_{\pi}$, the unique positive solution to (106). To that end, we recall the characterization in (128). Using the claim that $G_{e}(\pi)$ and $\widehat{\mathcal{G}}_{e}(\pi)$ are asymptotically equivalent (in particular, in the sense of (105)), we have
$\widehat{\chi}_{\mathrm{ave}}\simeq\frac{1}{d^{2}}\operatorname{tr}\left(\left[\widehat{\mathcal{G}}_{e}(\pi)\right]_{\setminus 0}\cdot\left[I\otimes E_{\mathrm{tr}}\right]\right).$ (133)
To compute the first term on the right-hand side of the above estimate, we directly invert the block matrix $\widehat{\mathcal{G}}_{e}(\pi)$ in (130). Recall that $\Pi_{e}$ is chosen in the forms of (85) and (92). It is then straightforward to verify that
$\widehat{\mathcal{G}}_{e}=\begin{bmatrix}\bar{c}&-\bar{c}\,{\bar{q}}^{\top}\\ -\bar{c}\,\bar{q}&I\otimes F_{E}(\widehat{\chi}_{\mathrm{ave}})+\bar{c}\,\bar{q}{\bar{q}}^{\top}\end{bmatrix}$, (134)
@@
$F_{E}(\chi)=\Big(\frac{\tau}{1+\chi}E_{\mathrm{tr}}+\pi B+\lambda\tau I_{d+1}\Big)^{-1},$ (135)
$\bar{q}=\frac{\tau}{(1+\widehat{\chi}_{\mathrm{ave}})\sqrt{d}}\operatorname{vec}\left(\begin{bmatrix}R_{\mathrm{tr}}&(1+\rho)b_{\mathrm{tr}}\end{bmatrix}F_{E}(\widehat{\chi}_{\mathrm{ave}})\right),$ (136)